The short-run marginal and average variable cost curves for a competitive firm are given by MC = 2+ 2q and AVC = 2 + q, respectively. The profit-maximizing level of output for a firm (q) is 4 and its total fixed cost (TFC) is $18. Which of the following must be true about the firm?

a) The firm is charging a price of $4 and covering its average variable cost, hence it should continue operating in the short-run.
b) The firm is charging $10 and will remain in the industry in the short run; but it is not covering its total costs and will consider leaving the industry in the long run.
c) The firm is charging a price of $4 and making a short-run loss, hence it should shut down immediately.
d) The firm is charging a price of $15 and making a zero profit, hence it should shut down eventually.
e) The firm is charging a price of $16 and making a negative profit, hence it should exit the industry in the long-run.
1
Expert's answer
2015-04-27T09:18:17-0400
MC = 2+ 2q, AVC = 2 + q, q = 4, TFC = $18.
AVC = 2 + 4 = 6. If P = 4, then P < AVC and firm must shut down, because it can't cover even its variable costs.
So, the right answer is:
c) The firm is charging a price of $4 and making a short-run loss, hence it should shut down immediately.